# VergeIO: Depth-Aware Eye Interaction on Glasses

### v1

@@ 5.5. Generalizing to virtual depth cues
To test, this, we developed a prototype application on the Meta Quest 2 [60], replicating the Brock string setup described in Sec.4 within a Unity-based VR environment. Beads were virtually placed at 30, 70, and 200 cm from the user’s eye center (Fig. 16a). We then invited a single participant (P4) to perform the same ten rounds of vergence gestures following the same procedure described in Sec. 4.
Using the frozen model trained on the Brock string dataset on all participants, we achieved an accuracy of 88.33% and 100% for six and four vergence-gestures classification respectively (Fig. 16c). This result shows that (1) virtual depth overlays can reliably induce vergence, even though they are rendered on a fixed display, and (2) a model trained on real-world depth cues can successfully generalize to virtual depth cues.
Figure 16. Experimental setup to evaluate generalization to (a) virtual depth cues and (b) approximate depth cues. Results are shown as confusion matrix as (c) and (d) respectively.
@@ 5.6. Generalizing to approximate depth cues
We next consider if our system can generalize to approximate physical depth cues with the user’s thumbs and a distant physical reference point (Fig. 16b). Specifically, the user fully extends one arm, rests the other hand at the elbow of the extended arm, and aligns both thumbs with a far-field object. This setup approximates the precise distances used to train the system where 30 and 70 cm, correspond approximately to half and full arm’s length for an average adult [61]. The exact depth of the far-field cue is less critical, as vergence-induced EOG signal changes diminish with increasing distance (Fig. 3).
@@ 5.7. Motion artifact and noise removal performance
We evaluated our motion artifact and noise removal pipeline using a dataset collected from a single participant (Table 2). To avoid data leakage, we ensured that no segments from the same recording appeared in both the training and validation sets. The model achieved an average accuracy of $0.986\pm 0.009$ (Fig. 14c).
Our results show that the model correctly classified 472 of 480 noise segments, with the remaining 8 misclassified as vergence. These false positives primarily consisted of standing and saccadic movements. Below in Sec. 5.8, we demonstrate that such misclassifications can be effectively mitigated by introducing a distinct preamble to activate and deactivate the system, and thus would not pose a practical issue. Conversely, 0 of 102 vergence segments were misclassified as noise, this is because all the vergence movements exceed the amplitude threshold used to classify noise.
@@ 5.8. False positives from non-vergence activity
Using our frozen classifier trained on data described in Sec.3.4 we obtain a false positive rate across conditions rangeing from 0.43% to 5.49% (Fig. 14d). To reduce the false positive rate, we use a preamble gesture, raising the brow, as a means of activating and deactivating the system. The motion is distinct from other movements, and results in an EOG amplitude that is significantly higher than vergence (Fig. 9), and it is easy and quick to perform. After the preamble, the false positive rate reduces to 0% for all conditions. These results show that by using a brow-raise as a preamble gesture, VergeIO can significantly reduce the false positive rate.
@@ 5.9. Power consumption
We measured the power consumption of our system using the Monsoon High Voltage Power Monitor [62] at 3.7 V, over 5 runs and averaged them, with each run lasting 1 minute at the lowest clock speed of 24 MHz.
The power consumption of the two amplifier front-ends was 3.3 $\pm$ 0.004 mW. This is lower than other eye tracking systems, in particular mmWave radios [8], and microphones and speakers [9], though we note that our system is purpose-built for the specific task of vergence gesture sensing not eye-tracking.
@@
With a 570 mAh rechargeable lithium-ion battery as in Google Glass [63], the system would operate for approximately 197 hours (8.2 days). This power draw is effectively negligible, as the overall battery life of Google Glass is dominated by other processes that resulting in a typical battery life of 8 hours [53].
@@ 6. Example applications
Vergence-driven varifocal glasses. We present a prototype of a mechanically tunable varifocal glass system. These glasses are useful for individuals with presbyopia, an age-related vision condition that affects more than a billion people [64] that reduces the eye’s ability to focus at different distances [65]. Importantly, while presbyopia impairs the focusing ability of the eye lens, individuals can still perform vergence movements between near and far distances [66].
Our prototype (Fig. 17a) instruments an existing pair of varifocal glasses [67] with a small servo motor [68] ($22.8\,\mathrm{mm}\times 12.2\,\mathrm{mm}\times 28.5\,\mathrm{mm}$) which is coupled to a knob to adjust the focal length. The motor weighs 29 g, which is relatively modest compared to commercial eye-tracking glasses such as the Tobii Pro Glasses 3, which weigh 388 g [14].
To show that our adjustable varifocal glasses can support focusing at distances of 30, 70, and 200 cm, we conducted a laser-based experiment to measure the effective focal lengths of the lens. The goal was to determine the focal length corresponding to different rotations of the glasses’ adjustment knob. In the setup (Fig. 17b), a laser pointer was aimed at a distant wall with the varifocal lens placed in between. Let $d$ be the distance from the laser to the lens, and $d^{\prime}$ the distance from the lens to the wall. We adjusted $d^{\prime}$ by moving the varifocal glasses until the laser spot appeared in focus, then computed the effective focal length $f$ using the lens equation [69]: $f=\frac{d\cdot d^{\prime}}{d+d^{\prime}}$. In our setup, the total distance between the laser and the wall was 12 m, which allowed us to measure focal lengths up to 3 m.
[…]
Stereoscopic depth rendering of virtual objects. Modern AR displays often render virtual objects at varying depths using stereoscopic rendering, which displays slightly different images to each eye to create a sense of depth from binocular disparity [70, 71]. Vergence sensing can be used to align virtual content with the user’s depth of focus. Importantly, although all virtual content is rendered on a screen located at a fixed physical distance, stereoscopic rendering can still induce vergence.
@@
Fig. 18 illustrates the stereoscopic depth estimation formula [72]: $D=d\cdot\left(\frac{L}{e}-1\right)$ which computes the required separation $D$ between the left and right projected images to make an object appear at a desired virtual depth $e$. Here, $L$ is the distance between the eyes and the screen, and $d$ is the interpupillary distance.
In the example, the user is positioned $L=200~{}cm$ from the screen, with an interpupillary distance of $d=5~{}cm$. To render an object at $e=30~{}cm$ and $e=70~{}cm$, the images would need to be separated by $D=28.3~{}cm$ and $D=9.3~{}cm$ respectively. While to render it at $e=200~{}cm$, which matches the physical distance to the screen, the images should overlap, resulting in zero separation.
@@ 7. Discussion
Learnability of vergence. Vergence movements have been shown to be a learnable and controllable skill even without visual reference points, enabling users to perform intentional depth-based eye gestures with training in a few minutes [73, 74, 4, 3, 75, 42]. An example voluntary vergence control is perceiving 3D autostereogram [76] using cross-eyed or wall-eyed viewing.
@@
Eye tracking using EOG signals. While the EOG signal can estimate horizontal and vertical eye movements based on amplitude changes, it is not well suited for precise gaze tracking due to its low spatial resolution. Additionally, it captures relative changes in eye position which can lead to error accumulation over time.
@@
However, we note that eye tracking is not an essential feature for commercial smart glasses, with devices like the Amazon Echo Frames [12] relying solely on speakers and microphones for interaction, to avoid the added cost and power demands of eye-tracking cameras.
@@
Vergence duration. Vergence is a naturally brief process and does not require prolonged fixation by users on the close or far object. In our user study, across all participants, the EOG signal from vergence had a duration of 0.737s ± 0.170 s, without requiring extended dwell periods. This distribution suggests that vergence can enable low-latency and responsive interaction across different users.

### v2

@@ 5.5. Generalizing to virtual depth cues
To test this, we developed a prototype application on the Meta Quest 2 (met, 2026), replicating the Brock string setup described in Sec. 4 within a Unity-based VR environment. Beads were virtually placed at 30, 70, and 200 cm from the user’s eye center (Fig. 16a). We then invited a single participant (P4) to perform the same ten rounds of vergence gestures following the same procedure described in Sec. 4.
Using the frozen model trained on the Brock string dataset on all participants, we achieved an accuracy of 88.33% and 100% for six and four vergence-gestures classification respectively (Fig. 16b). This result shows that (1) virtual depth overlays can reliably induce vergence, even though they are rendered on a fixed display, and (2) a model trained on real-world depth cues can successfully generalize to virtual depth cues.
Figure 16. Experimental setup to evaluate generalization to (a) virtual depth cues and (c) approximate depth cues. Results are shown as confusion matrix as (b) and (d) respectively.
@@ 5.6. Generalizing to approximate depth cues
We next consider if our system can generalize to approximate physical depth cues with the user’s thumbs and a distant physical reference point (Fig. 16c). Specifically, the user fully extends one arm, rests the other hand at the elbow of the extended arm, and aligns both thumbs with a far-field object. This setup approximates the precise distances used to train the system where 30 and 70 cm, correspond approximately to half and full arm’s length for an average adult (Roberts et al., 1959). The exact depth of the far-field cue is less critical, as vergence-induced EOG signal changes diminish with increasing distance (Fig. 3).
@@ 5.7. Motion artifact and noise removal performance
We evaluated our motion artifact and noise removal pipeline using a dataset collected from a single participant (Table 2). To avoid data leakage, we ensured that no segments from the same recording appeared in both the training and validation sets. The model achieved an average accuracy of $0.986\pm 0.009$ (Fig. 13b).
Our pipeline demonstrated robustness by correctly classifying 472 out of 480 noise segments, while the remaining 8 instances were misclassified as vergence. These false positives primarily occurred during standing or saccadic movements. As discussed in Sec. 5.8, such misclassifications are effectively mitigated through a distinct preamble for system activation, ensuring they do not present practical usability issues. In contrast, none of the 102 vergence segments was misclassified as noise because all vergence movements consistently exceeded the amplitude threshold defined for noise classification.
@@ 5.8. False positives from non-vergence activity
Using our frozen classifier trained on data described in Sec. 3.4 we obtain a false positive rate across conditions ranging from 0.43% to 5.49% (Fig. 13c). To reduce the false positive rate, we use a preamble gesture, raising the brow, as a means of activating and deactivating the system. The motion is distinct from other movements, and results in an EOG amplitude that is significantly higher than vergence (Fig. 8), and it is easy and quick to perform. After the preamble, the false positive rate reduces to 0% for all conditions. These results show that by using a brow-raise as a preamble gesture, VergeIO can significantly reduce the false positive rate.
@@ 5.9. Power consumption
We measured the power consumption of our system using the Monsoon High Voltage Power Monitor (mon, 2026) at 3.7 V, over 5 runs and averaged them, with each run lasting 1 minute at the lowest clock speed of 24 MHz.
The power consumption of the two amplifier front-ends was 3.3 $\pm$ 0.004 mW. This is lower than other eye tracking systems, in particular mmWave radios (Ma et al., 2025), and microphones and speakers (Li et al., 2024), though we note that our system is purpose-built for the specific task of vergence gesture sensing not eye-tracking.
@@
With a 570 mAh rechargeable lithium-ion battery as in Google Glass (ggl, 2026a), the system would operate for approximately 197 hours (8.2 days). This power draw is effectively negligible, as the overall battery life of Google Glass is dominated by other processes that resulting in a typical battery life of 8 hours (ggl, 2026b).
@@ 5.10. System latency
5.11. User Experience Survey
We evaluated the user experience during the experiment with a survey. We asked the following questions:
Figure 17. Demographic summary of participants in the user study. Survey data was collected on $n=20$ participants of the data collection procedure.
Figure 18. User experience survey ($n=20$). Histograms summarize user perceptions across (a) ease of maintaining electrode contact, (b) ease of facial muscle movement while wearing the prototype, (c) concern regarding accidental gestures triggers, (d) interaction-induced fatigue, and the social acceptability and comfort of the interaction in (e) private and (f) public settings.
  1. (1)
    How easy is it to keep the electrodes on? (1 = Hard, 5 = Very easy to keep on)
  2. (2)
    How easy was it to move your facial muscles (e.g. raising your eyebrows, closing your eyelids)? (1 = Very difficult, 5 = Not Difficult)
  3. (3)
    How concerned are you about accidentally performing these eye gestures in your daily life? (1 = Very concerned, 5 = Not at all)
  4. (4)
    To what extent did you feel physical or mental fatigue when performing an eye gesture with the system? (1 = High fatigue, 5 = No fatigue)
  5. (5)
    How comfortable would they feel performing these eye gestures in a public setting? (1 = Very uncomfortable, 5 = Very comfortable)
  6. (6)
    How comfortable would you feel performing these eye gestures in a private setting? (1 = Very uncomfortable, 5 = Very comfortable)
The survey results are summarized in Fig. 18. Participants ($n=20$) had an average age of $25.75\pm 11.1$ years and height of $171.83\pm 7.65$ cm. The system demonstrated reliable wearability, as participants reported that the electrodes remained securely in place with an average comfort score of $3.85\pm 0.93$ out of 5. Participants found the system did not impede natural expressions, with an average score of $4.10\pm 0.98$ for the ease of performing facial movements such as raising eyebrows, closing eyelids, or speaking. These findings indicate that VergeIO maintains stable electrode contact across users while ensuring that facial actions can be performed without interference.
In terms of social acceptability, the average comfort rating reached $4.75\pm 0.55$ in private settings but decreased to $3.75\pm 0.91$ in public environments. Despite a slight decrease in comfort while in public, the scores suggest that users are generally willing to incorporate the device into their daily routines. The concern about accidentally performing the gesture in daily life received an average score of $3.75\pm 1.25$, indicating that users are confident in performing the movement intentionally on command. Finally, participants reported a fatigue level of $3.55\pm 0.89$, where a score of 5 represents a total absence of fatigue. This result implies that the interaction remains feasible for frequent use throughout the day, which supports the long-term viability of continuous vergence monitoring.
@@ 6. Example applications
6.1. Vergence-driven varifocal glasses.
We present a prototype of a close-loop, mechanically tunable varifocal glass system. This design aligns to a growing interest in both academia and industry on ”autofocals”, which aims to create intelligent eyewear that re-enables dynamic focusing (Padmanaban et al., 2018). Such technology is designed to address a wide spectrum of accommodation-related vision issues, including post-cataract surgery recovery, pediatric amblyopia and presbyopia, the most widespread vision impairment that affects more than a billion people worldwide (Holden et al., 2008). Recent commercial entries, such as IXI (ixi, 2026), ViXion01 (ViX, 2026), and Elcyo (elc, 2026), have demonstrated the market’s appetite for active focal adjustment. These solutions rely on vision-based modalities, which can be power-intensive for continuous wearable use.
In contrast, we present the first attempt to leverage EOG signals as a low-power, lightweight modality for automatic varifocal control. Importantly, although the eye lens’ focusing ability is impaired, individuals can still perform vergence movements betweeen near and far distances  (Padmanaban et al., 2018). We have also validated this in Sec. 5.2, where our system achieved $77.90\pm 1.77\%$ 6-gesture and $94.38\pm 0.62\%$ 4-gesture accuracy for individuals with presbyopia.
Our prototype (Fig. 19a) instruments an existing pair of varifocal glasses (var, 2026) with a small servo motor (ser, 2026) ($22.8\,\mathrm{mm}\times 12.2\,\mathrm{mm}\times 28.5\,\mathrm{mm}$) which is coupled to a knob to adjust the focal length. The motor weighs 29 g, which is relatively modest compared to commercial eye-tracking glasses such as the Tobii Pro Glasses 3 (388 g) (tob, 2026).
To show that our adjustable varifocal glasses can support focusing at distances of 30, 70, and 200 cm, we conducted a laser-based experiment to measure the effective focal lengths of the lens. The goal was to determine the focal length corresponding to different rotations of the glasses’ adjustment knob. In the setup (Fig. 19b), a laser pointer was aimed at a distant wall with the varifocal lens placed in between. Let $d$ be the distance from the laser to the lens and $d^{\prime}$ the distance from the lens to the wall. We adjusted $d^{\prime}$ by moving the varifocal glasses until the laser spot appeared in focus, then computed the effective focal length $f$ using the lens equation (len, 2026): $f=\frac{d\cdot d^{\prime}}{d+d^{\prime}}$. The total distance between the laser and the wall was 12 m, which allowed us to measure focal lengths up to 3 m.
This procedure was repeated for five normalized knob rotations ranging from 0.0 to 1.0. As shown in Fig. 19c, the effective focal length spans approximately $25.58\,\text{cm}$ to over $300\,\text{cm}$, covering the key distances required for near-to-far visual accommodation in individuals with presbyopia.
6.2. Stereoscopic depth rendering of virtual objects.
Modern AR displays often render virtual objects at varying depths using stereoscopic rendering, which displays slightly different images to each eye to create a sense of depth from binocular disparity (Blondé et al., 2010; Wan et al., 2004). Vergence sensing can be used to align virtual content with the user’s depth of focus. Importantly, although all virtual content is rendered on a screen located at a fixed physical distance, stereoscopic rendering can still induce vergence.
@@
Fig. 20 illustrates the stereoscopic depth estimation formula (Rule, 1941): $D=d\cdot\left(\frac{L}{e}-1\right)$ which computes the required separation $D$ between the left and right projected images to make an object appear at a desired virtual depth $e$. Here, $L$ is the distance between the eyes and the screen, and $d$ is the interpupillary distance.
In the example, the user is positioned L=200cmL=200~cm from the screen, with an interpupillary distance of d=5cmd=5~cm. To render an object at e=30cme=30~cm and e=70cme=70~cm, the images would need to be separated by D=28.3cmD=28.3~cm and D=9.3cmD=9.3~cm respectively. While to render it at e=200cme=200~cm, which matches the physical distance to the screen, the images should overlap, resulting in zero separation.
@@ 7.1. Learnability of vergence
Vergence have been shown to be a learnable and controllable skill even without visual reference points, enabling users to perform intentional depth-based eye gestures with training in a few minutes (McLin and Schor, 1988; Hamori et al., 1982; Zhang et al., 2024; Kirst and Bulling, 2016; Kudo et al., 2013; Ruan et al., 2018). An example voluntary vergence control is perceiving 3D autostereogram (Tyler and Clarke, 1990) using cross-eyed or wall-eyed viewing.
7.2. Eye tracking using EOG signals
@@
While the EOG signal can estimate horizontal and vertical eye movements based on amplitude changes, it is not well suited for precise gaze tracking due to its low spatial resolution. As noted in prior literature, the spatial resolution of EOG is typically limited to approximately 1.5° (Young and Sheena, 1975). Additionally, it captures relative changes in eye position which can lead to error accumulation over time.
@@
However, we note that eye tracking is not an essential feature for commercial smart glasses, with devices like the Amazon Echo Frames (ama, 2026) relying solely on speakers and microphones for interaction, to avoid the added cost and power demands of eye-tracking cameras.
7.3. Vergence duration
@@
Vergence is a naturally brief process and does not require prolonged fixation by users on the close or far object. In our user study, across all participants, the EOG signal from vergence had a duration of 0.737s ± 0.170 s, without requiring extended dwell periods. This distribution suggests that vergence can enable low-latency and responsive interaction across different users.
7.4. Anthropometric adaptability of VergeIO
A fundamental challenge for EOG-based eye movement sensing systems is balancing robust electrode-skin contact with user comfort. Our design addresses the challenge by employing the nose-bridge as a natural contact point, integrated with flexible sliding arms and a curved mastoid structure. User evaluations in Sec. 5.11 confirm that the design effectively maintains stable electrode contact without compromising comfort. Beyond passive mechanical optimization, future sensing systems could implement closed-loop adjustment mechanisms driven by real-time signal quality and user experience. While medical-grade wearables like emteqPRO (Gnacek et al., 2022) rely on EMG-based impedance for contact validation, future EOG systems on smart glasses could similarly leverage signal quality indices and user feedback to dynamically guide device positioning or trigger automated micro-adjustments.
7.5. Limitations and Future Work
Robustness against remounting and cross-user variation. While VergeIO expands detectable eye movements on EOG-based glasses, ensuring robustness across usage contexts remains a challenge. Our evaluation indicates reduced accuracy in cross-session ($77.78\%$) and the cross-user ($77.36\%$) evaluations, attributed to the signal drift from electrode placement changes and physiological variability between subjects. Future work will leverage larger-scale datasets and advanced deep learning techniques, such as domain adaptation. This approach aims to learn subject-invariant representations, mitigating deviations caused by remounting and individual differences.
Preamble mechanism to mitigate accidental triggers. To ensure system reliability against daily facial activities (e.g. chewing, speaking), our current design incorporates raising eyebrows as an explicit preamble for activation. Although this mechanism effectively mitigates false positives, requiring a deliberate physical gesture may be intrusive in social contexts and interrupt natural interaction flows. We aim to evolve towards implicit activation mechanisms in future iterations, potentially by fusing EOG with lightweight IMU data or leveraging context-aware algorithms to seamlessly discern interaction intent without conscious user effort.
Effectiveness in long-term usage. As our evaluation was conducted primarily in controlled settings, future longitudinal in-the-wild studies are essential to validate the system’s resilience to dynamic physiological factors, such as perspiration and skin impedance changes, during prolonged daily usage.
8. Conclusion
This paper presented VergeIO, the first EOG-based glasses system that enables robust, depth-aware eye interaction by sensing vergence. With an optimized electrode layout with positive electrodes in the temples and a shared negative electrode in the nose bridge, we overcome the limitations of previous designs that were unable to distinguish asymmetric electrical responses of vergence. Our evaluation across 20 users demonstrated classification accuracies of 83–98% for personalized models and 77–97% for cross-user generalization without calibration. With a front-end power consumption of only 3 mW and a real-time processing latency of 15 ms, the system is designed for continuous, always-on use. We envision that the sensing capability introduced by VergeIO will broaden the scope of eye interaction available on EOG-based glasses, serving as a lightweight and privacy-preserving primitive that enables a new generation of depth-aware applications in consumer electronics, healthcare, and safety-monitoring domains.
9. LLM Use Clarification
Gemini-3 Pro was used for generating human characters in Fig. 9, as well as for grammar correction and text refinement throughout this paper.